# ref.bib **footnotetext: Equal contribution

### v1

@@ 3.3.1 Frame Annotation Module (FAM)
As described in Sec. 3.2, we first collect raw videos from existing VAD datasets like CUHK Avenue [cuhk-avenue], UCF-Crime [ucf-crime], UCSD-Ped [ucsd-ped], and also add additional challenging AI generated videos to the mix. For the VAD datasets, the bounding box annotations were already provided for a subset of the videos from these datasets. Thus, we only annotate the anomalies present in the AI-generated videos. In this stage, we first break down the raw videos into their constituent image frames. Second, we select and filter 10 consecutive frames from the video, which contains the inconsistency. We annotate these selected frames with a bounding box mentioning the type of the inconsistency. We consider the following inconsistency types: ‘Sudden Appearance’, ‘Unnatural Transform’, ‘Disappearance’, ‘Pass-through, and ‘Unnatural Appearance’. Fig. 4 shows the annotated ‘Unnatural Transform’ inconsistency affecting the kangaroo’s legs and tails.
@@ 3.3.2 Caption Generation Module (CGM)
The second stage of our data collection process involves the Caption Generation Module (CGM), which uses the annotated video frames from FAM, to generate a high quality and detailed caption which describes the inconsistency, along with the general events in the video. To generate the caption, we design a specialised custom prompt, and use the recently released GPT-4o [gpt4-o] LMM, which has shown both impressive performance gains and cost savings. Thus, GPT-4o model takes in our custom prompt, along with the annotated frames to generate the descriptive video caption as shown in Fig. 4.
@@ 3.3.3 Question Answer Generation Module (QAGM)
The final stage of our VANE-Bench construction process involves using the generated caption from CGM, and the annotated frames from FAM to output the final high-quality, and challenging Question and Answer (QA) pairs. We create another custom prompt which we pass to the GPT-4o model, along with caption, and the annotated frames as input to generate the QA pairs. The selected raw frames containing the inconsistency, and their corresponding generated QA pairs form our VANE-Bench dataset.
@@ 4 Experiments and Results
Video-LMMs. We evaluate the anomaly detection and comprehension capabilities of both open-source and closed-source models. Among the open-source models, we evaluate 7 recent Video-LMMs, including Video-LLaVA [lin2023video], TimeChat [ren2023timechat], MovieChat [song2023moviechat], LLaMA-ViD [li2023llama], VideoChat [li2023videochat], Video-ChatGPT [maaz2023video], and Video-LLaMA-2 [zhang2023video]. For evaluating closed-source models, we use Gemini-1.5 Pro [Gemini] and GPT-4o [gpt4-o].
Evaluation Protocol. For the evaluation of Gemini and GPT-4o, we utilize their respective official APIs, with each model receiving 10 video frames as input. The 10 frames are selected in a manner that encompasses all or the majority of the inconsistencies present in the video. In cases where an anomaly spans a longer duration, we sample multiple sets of 10 frames to ensure comprehensive coverage. As GPT-4o does not inherently support videos, we input the video clips as 10 frames to the GPT API, accompanied by the corresponding Visual Question-Answering (VQA) query. For each model under assessment, we generate responses to the questions independently and without retaining the conversation history. Few models such as Moviechat, output hallucinated responses when instructed to answer the query. In such cases, we consider the hallucinated responses as incorrect answer, due to the inability of the model to comprehend the situation in the video.
Table 1: Evaluation results of Video-LMMs across different types of video samples on the VANE benchmark. We present results for both open-source and closed-source models. The first five rows show results on AI-generated videos and last four contains results on real world anomaly datasets.
@@ 4.1.1 Evaluation on Video-LLMs
AI-Generated anomalies. The AI-generated videos in our dataset are derived from five distinct models: SORA, OpenSORA, Runaway Gen-2, VideoLCM, and Modelscope-T2V. In the majority of these videos, the anomalies are subtle and not readily apparent, even to the human eye. As previously stated in section 3.2, the synthetic anomalies can manifest in five different forms. The performance of open-source models in detecting anomalies in these videos is subpar. Although closed-source models outperform their open-source counterparts, their overall comprehension and detection of anomalies in the videos remain inadequate. This indicates that even robust closed-source models encounter difficulties in identifying subtle anomalies within the videos.
Figure 5: Human vs Video-LMMs’ performance on SORA. Performance comparison of humans vs Video-LMMs on VQA task of detecting anomalies in SORA dataset. We find that closed-source Video-LMMs perform comparably to humans while open-source Video-LMMs struggle to detect the subtle anomalies.
Real-world anomalies. Our real-world anomaly datasets benchmark, as discussed in section 3.2, comprises four real-world datasets and focuses on detecting crime-related irregularities, atypical pedestrian behavior, and unusual events. These anomalies are prevalent in real-world scenarios. In our analysis, we find that open-source models encounter difficulties in locating and identifying these anomalies. As shown in the Table 1, these models perform poorly on these datasets. Conversely, we observe that closed-source models excel at detecting such real-world anomalies, indicating that they can effectively differentiate between the unusual events in the real-world scenarios. This can be attributed to the fact that these models are trained on a vast amount of existing real-world, internet-scale data.
@@ 4.1.2 Human Evaluation
We conducted a human evaluation on SORA-generated videos, which contain subtle and challenging anomalies that are difficult for humans to detect (see Fig. 1 top row) in a single viewing. Moreover, most of the video clips contain a multitude of foreground and background characters and elements, which makes it difficult for humans to focus on the inconsistencies within the short time frame. Some of the questions also specifically inquire about inconsistencies present in the background characters of the clips, rather than the foreground ones. To ensure fairness, our human evaluation was conducted under a set of rules, which include showing all 10 frames of the video to the human evaluator only once, followed by the question. Our human evaluation comparisons are presented in Fig 5. While humans outperform open-source models in detecting these subtle anomalies, their performance remains sub-optimal. This indicates that, with the advancements in video generation techniques, there is a pressing need for more sophisticated and effective Video-LMMs capable of assisting in the detection of such challenging cases capable of evading human eyes as well.
Figure 6: Inconsistency in Predictions: Left: Video-ChatGPT and VideoChat predict accurately, while Video-LLAMA selects incorrectly. Right: With a rephrased query, predictions shift. Video-ChatGPT and VideoChat err, whereas Video-LLAMA predicts correctly. This indicates the sensitivity of Video-LMMs towards the phrasing of query.
@@ 4.2 Additional Analysis
Inconsistencies in Predictions. We find that, in the majority of cases, open-source Video-LMMs generate different results when prompted to answer the same query multiple times. Fig. 6 illustrates a sample example where the same questions were posed twice to the corresponding Video-LMMs, yielding different responses. In some instances, the answers generated by the Video-LMMs in both rounds were dissimilar and incorrect. However, we also found cases where Video-LMMs initially produced the correct answer, followed by an incorrect answer to the same query, albeit phrased slightly differently. This suggests that the majority of these open-source Video-LMMs struggle to comprehend the same query when presented in a different manner, leading to inconsistent and paradoxical predictions. In contrast, closed-source Video-LMMs are less prone to such inconsistent predictions and consistently produce the same output for the same queries, regardless of how they are phrased, indicating a superior comprehension of language.
@@ 5 Conclusion
[…]
@@
  • Additional Findings and Qualitative Results (Appendix A)
@@ Appendix A Additional findings and qualitative results
A.1 Qualitative Results
@@
A.2 VANE-Bench frequent instances
Figure 7 presents a word cloud visualization, highlighting the most frequently occurring keywords within the correct option set of the VANE-Benchmark dataset. These prominent words are indicative of objects or human attributes in the videos that are most likely to exhibit anomalous behavior. From the figure, the most frequently occurring keyword is "Face" which indicated that the synthetically generated videos most likely struggle to generate a perfect human face.
@@ Appendix B Additional results on Prediction inconsistency
Figure 9: Prediction Inconsistency: Figure shows the response of Video-LMMs to the VQA task of detecting anomalies in the video. The correct answer is written in bold in the user query. We find that majority of Video-LMMs struggle to answer the questions correctly.
@@ Appendix C Importance of Frame annotation module
Since the video inconsistencies present in state of the art AI models like SORA are quite subtle, and hard to detect, our Frame Annotation Module (FAM) ensures that we are able to generate high-quality and accurate captions for these videos. As shown in Fig. 10, without FAM, the generated caption is not able to describe the sudden appearance of the kangaroo’s right foot near its tail. Further, the caption generated without our FAM is also not able to describe the extra set of paws which appear suddenly from the legs of the cat. Thus, FAM plays an important role in curating high-quality and accurate video captions.
@@ Appendix D Implementation Details
We use the official code of each open-source Video-LMM for evaluation. Each of these codes are implemented in pytorch framework. We evaluate each one of them on a single NVIDIA A100 40GB GPU. For closed source Video-LMMs, we use their respective API for evaluation. We use GPT-4o [gpt4-o] as our LMM to generate the captions and the final QA pairs in VANE-Bench. Next, we describe the prompts used in our Caption Generation Module (CGM), Question Answer Generation Module (QAGM), and in evaluating various Video-LMMs on VANE-Bench in the subsequent subsections.
@@ D.2 Question Answer Generation Module (QAGM)
Text Prompt: You are given a video input, which is generated by a state of the art AI algorithm. Thus, these videos look very natural and almost realistic, but they are actually synthetic and generated by an AI algorithm. The videos may have some inconsistencies or anomalies present in them, which are generally localized to only a specific location in the video as identified by the green bounding boxes in the video. The rest of the video appears completely natural or realistic. These specific inconsistency may last for only a few frames of the video, or may last for the entire video itself. The inconsistency or anomalies in the video are generally events and phenomenon which is not observed in real world and physical scenarios. You will also be given a caption as input which describes the video, along with the specific inconsistency present in the video. Based on the given video and caption input, your task is to formulate 3 diverse and misleading questions to test whether the multi-modal LLM model can correctly identify the options based on the inconsistencies present in the video or not. So, your generated questions should give the model few options to choose from to make its answer, and these options should be of high quality and also have misleading choices so that you can test deeper level of understanding of these multi-modal LLM models. Thus, the goal of these questions is to accurately assess the multi-modal LLM’s ability to accurately identify the inconsistencies present in the video. Generate questions that comprise both interrogative and declarative sentences, utilizing different language styles, and provide an explanation for each. Your response should be presented as a list of dictionary strings with keys ’Q’ for questions and ’A’ for the answer. Follow these rules while generating question and answers:
1. Do not provide answers in the question itself. For example, the ground-truth attribute or component which makes the video scene unusual should never be mentioned in the question itself.
@@
9. You must only follow the below output format, and strictly must not output any other extra information or text. Your output format should be strictly as follows, without any additional information or text:

### v2

@@ 3.3.1 Frame Annotation Module (FAM)
As described in Sec. 3.2, we first collect raw videos from existing VAD datasets like CUHK Avenue Lu et al. (2013), UCF-Crime Sultani et al. (2018), UCSD-Ped Li et al. (2014), and also add additional challenging AI generated videos to the mix. For the VAD datasets, the bounding box annotations were already provided for a subset of the videos from these datasets. Thus, we only annotate the anomalies present in the AI-generated videos. In this stage, we first break down the raw videos into their constituent image frames. Second, we select and filter 10 consecutive frames from the video that contain the inconsistency. We annotate these selected frames with a bounding box mentioning the type of inconsistency. We consider the following inconsistency types: ‘Sudden Appearance’, ‘Unnatural Transform’, ‘Disappearance’, ‘Pass-through, and ‘Unnatural Appearance’. Fig. 4 shows the annotated ‘Unnatural Transform’ inconsistency affecting the kangaroo’s legs and tails.
@@ 3.3.2 Caption Generation Module (CGM)
The second stage of our data collection process involves the Caption Generation Module (CGM), which uses the annotated video frames from FAM to generate a high-quality and detailed caption which describes the inconsistency, along with the general events in the video. To generate the caption, we design a specialised custom prompt (Sec. D.1), and use the recently released GPT-4o OpenAI (2024) LMM, which has shown both impressive performance gains and cost savings. Thus, GPT-4o model takes in our custom prompt, along with the annotated frames to generate the descriptive video caption as shown in Fig. 4.
@@ 3.3.3 Question Answer Generation Module (QAGM)
The final stage of our VANE-Bench construction process involves using the generated caption from CGM, and the annotated frames from FAM to output the final high-quality, and challenging Question and Answer (QA) pairs. We create another custom prompt (Sec. D.2) which we pass to the GPT-4o model, along with caption, and the annotated frames as input to generate the QA pairs. The selected raw frames containing the inconsistency, and their corresponding generated QA pairs form our VANE-Bench dataset.
@@ 4 Experiments and Results
Table 1: Evaluation results of Video-LMMs across different types of video samples on the VANE benchmark. We present results for both open-source and closed-source models. The first five rows show results on AI-generated videos and last four contain results on real world anomaly datasets.
Video-LMMs. We evaluate the anomaly detection and comprehension capabilities of both open-source and closed-source models. Among the open-source models, we evaluate 7 recent Video-LMMs, including Video-LLaVA Lin et al. (2023), TimeChat Ren et al. (2023), MovieChat Song et al. (2023), LLaMA-ViD Li et al. (2023c), VideoChat Li et al. (2023b), Video-ChatGPT Maaz et al. (2023), and Video-LLaMA-2 Zhang et al. (2023a). For evaluating closed-source models, we use Gemini-1.5 Pro Google (2023) and GPT-4o OpenAI (2024).
Evaluation Protocol. For the evaluation of Gemini and GPT-4o, we utilize their respective official APIs, with each model receiving 10 video frames as input. The 10 frames are selected in a manner that encompasses all or the majority of the inconsistencies present in the video. In cases where an anomaly spans a longer duration, we sample multiple sets of 10 frames to ensure comprehensive coverage. As GPT-4o does not inherently support videos, we input the video clips as 10 frames to the GPT API, accompanied by the corresponding Visual Question-Answering (VQA) query. For each model under assessment, we generate responses to the questions independently and without retaining the conversation history. Few models, such as Moviechat, output hallucinated responses when instructed to answer the query. In such cases, we consider the hallucinated responses as incorrect answers due to the inability of the model to comprehend the situation in the video.
@@ 4.1.1 Evaluation on Video-LLMs
AI-Generated anomalies. The AI-generated videos in our dataset are derived from five distinct models: SORA, OpenSORA, Runaway Gen-2, VideoLCM, and Modelscope-T2V. In the majority of these videos, the anomalies are subtle and not readily apparent, even to the human eye. As previously stated in section 3.2, the synthetic anomalies can manifest in five different forms. As shown in Table 1, the performance of open-source models in detecting anomalies in these videos is subpar. Although closed-source models outperform their open-source counterparts, their overall comprehension and detection of anomalies in the videos remain inadequate. This indicates that even robust closed-source models encounter difficulties in identifying subtle anomalies within the videos.
Real-world anomalies. Our real-world anomaly datasets benchmark, as discussed in section 3.2, comprises four real-world datasets and focuses on detecting crime-related irregularities, atypical pedestrian behavior, and unusual events. These anomalies are prevalent in real-world scenarios. In our analysis, we find that open-source models encounter difficulties in locating and identifying these anomalies. As shown in Table 1, these models perform poorly on these datasets. Conversely, we observe that closed-source models excel at detecting such real-world anomalies, indicating that they can effectively differentiate between unusual events in real-world scenarios. This can be attributed to the fact that these models are trained on a vast amount of existing real-world, internet-scale data.
We provide results on additional latest Video-LMMs in Section A.1 of Supplementary.
Figure 5: Human vs Video-LMMs’ performance on SORA. Performance comparison of humans vs Video-LMMs on VQA task of detecting anomalies in SORA dataset. We find that closed-source Video-LMMs perform comparably to humans while open-source Video-LMMs struggle to detect subtle anomalies.
Figure 6: Inconsistency in Predictions: Left: Video-ChatGPT and VideoChat predict accurately, while Video-LLAMA selects incorrectly. Right: With a rephrased query, predictions shift. Video-ChatGPT and VideoChat err, whereas Video-LLAMA predicts correctly. This indicates the sensitivity of Video-LMMs towards query rephrasing.
@@ 4.1.2 Human Evaluation
We conducted a human evaluation on SORA-generated videos, which contain subtle and challenging anomalies that are difficult for humans to detect (see Fig. 1 top row) in a single viewing. Moreover, most of the video clips contain a multitude of foreground and background characters and elements, which makes it difficult for humans to focus on the inconsistencies within the short time frame. Some of the questions also specifically inquire about inconsistencies present in the background characters of the clips rather than the foreground ones. To ensure fairness, our human evaluation was conducted under a set of rules, which include showing all 10 frames of the video to the human evaluator only once, followed by the question. Our human evaluation comparisons are presented in Fig 5. While humans outperform open-source models in detecting these subtle anomalies, their performance remains sub-optimal. This indicates that, with the advancements in video generation techniques, there is a pressing need for more sophisticated and effective Video-LMMs capable of assisting in the detection of such challenging cases capable of evading human eyes as well.
@@ 4.2 Additional Analysis
Inconsistencies in Predictions. We find that, in the majority of cases, open-source Video-LMMs generate different results when prompted to answer the same query multiple times. Fig. 6 illustrates a sample example where the same questions were posed twice to the corresponding Video-LMMs, yielding different responses. In some instances, the answers generated by the Video-LMMs in both rounds were dissimilar and incorrect. However, we also found cases where Video-LMMs initially produced the correct answer, followed by an incorrect answer to the same query, albeit phrased slightly differently. This suggests that the majority of these open-source Video-LMMs struggle to comprehend the same query when presented in a different manner, leading to inconsistent and paradoxical predictions. In contrast, closed-source Video-LMMs are less prone to such inconsistent predictions and consistently produce the same output for the same queries, regardless of how they are phrased, indicating a superior comprehension of language. Refer to supplementary Section B for additional results.
@@
We provide more insights and discussions in Section A.4 of Supplementary material.
@@ 5 Conclusion
6 Limitations
Our VANE-Bench is the first benchmark for evaluating Video-LMMs on anomalous videos from both AI-generated and real-world sources. While we have done our best to ensure a high-quality evaluation of these Video-LMMs, certain limitations still manifest.
Our Question-answer pairs are designed to have 4 options. We design the instruct prompt to ensure that each Video-LMM outputs one out of 4 options. However, in some instances, the model outputs a hallucinated response and does not follow the instructions. As a result, we employ a post-response human-based filtration process, which involves an exhaustive verification and rectification of these errors. In our current setup, we mark these cases as wrong. We believe that future Video-LLMs will be more aligned with human intent and will follow human instructions appropriately.
Additionally, the video samples from the SORA are limited in VANE-Bench. This is due to the fact that SORA model is not open-source yet, hence we rely on publicly available samples of SORA for evaluation.
@@ 6 Limitations
  • Additional Findings and Results (Appendix A)
@@ Appendix A Additional Findings and Qualitative Results
A.1 Additional Quantitative Results
Video-LMMs are fewer in number compared to image-based multi-modal models, which limits the range of options available for evaluation. Given this scarcity, we selected 7 open-source and 2 closed-source LMMs that are currently among the most widely used. To ensure that our benchmark remains representative, we have also included additional results from other latest open-source LMMs: Ataallah et al. (2024a, b); Li et al. (2024a), as shown in Table 2. Our findings reveal that open-source models still lag behind their closed-source counterparts in performance, indicating that simply adding more models wouldn’t necessarily improve the overall representativeness of our benchmark. Our selected set, which includes both open-source and closed-source models, is already comprehensive, featuring state-of-the-art models like GPT-4o and Gemini-1.5 Pro. Given the limited number of Video-LMMs available, our reliance on this specific set of models is justified, as it accurately represents the current landscape of Video-LMM capabilities.
Benchmark Category LLaVA-NeXT MiniGPT4-Video Goldfish
SORA 11.59 10.74 26.47
OpenSORA 18.00 28.00 22.00
Runway Gen2 16.00 4.00 12.00
VideoLCM 10.57 17.64 18.26
Modelscope-T2V 10.41 20.83 16.66
UCFCrime 9.47 11.57 31.57
UCSD-Ped1 16.66 13.33 40.00
UCSD-Ped2 5.55 13.88 19.44
Table 2: Evaluation results of additional latest Video-LMMs across different types of video samples on the VANE benchmark. We present results for both open-source and closed-source models. The first five rows show results on AI-generated videos and last three contain results on real world anomaly datasets.
A.2 Qualitative Results
@@
A.3 VANE-Bench frequent instances
@@
Figure 7 presents a word cloud visualization, highlighting the most frequently occurring keywords within the correct option set of the VANE-Benchmark dataset. These prominent words are indicative of objects or human attributes in the videos that are most likely to exhibit anomalous behavior. From the figure, the most frequently occurring keyword is "Face" which indicates that the synthetically generated videos most likely struggle to generate a perfect human face.
@@
A.4 Additional Discussions on Experimental Results
Per anomaly performance: To give further insights, Figure 2 (left) of the main paper illustrates the performance of LMMs on each type of anomaly present in the AI-generated videos. We can observe that closed-source models like GPT-4o and Gemini-1.5 Pro consistently exhibit strong performance across all five anomaly categories compared to their open-source counterparts. This likely stems from their access to significantly larger training datasets and model parameters, allowing for a more robust understanding of visual anomalies. Conversely, open-source models exhibit fluctuating performance depending on the anomaly type. We also note that open-source models struggle, especially with the “disappearance” anomaly. We believe that it might be because of the fact that these models are trained on datasets focusing on the presence of objects and actions, and hence being more biased towards presence. Further, we believe that open-source models suffers from limited temporal reasoning capability and often use short-term mechanisms that limit their ability to track objects over time. The lack of datasets focusing on anomalies like “disappearance” also limits the model’s capability to detect such patterns.
Higher performance of some LMMs: As seen in Table 1 and Figure 5 of main paper, we notice that some open-source LMMs perform better than their counterparts. For instance, we notice Video-ChatGPT achieves higher performance compared to other open-source models. We believe that it might be because of the following two reasons: 1. Training Data: While most open-source models rely solely on web-scraped video captioning data, Video-ChatGPT incorporates a manually annotated video instruction dataset specifically designed for video understanding. This provides the model with a more direct and targeted learning experience, potentially enhancing its sensitivity to anomalies. 2. Two-Stage Training: Video-ChatGPT employs a two-stage training process involving both video-language pre-training and instruction tuning. This enables the model to first develop a strong understanding of general video semantics and then refine its ability to follow user instructions and reason about specific events within videos.
@@ Appendix B Additional results on Prediction inconsistency
Figure 9: Prediction Inconsistency: Figure shows the response of Video-LMMs to the VQA task of detecting anomalies in the video. The correct answer is written in bold in the user query. We find that the majority of Video-LMMs struggle to answer the questions correctly.
@@ Appendix C Importance of Frame annotation module
Since the video inconsistencies present in state of the art AI models like SORA are quite subtle, and hard to detect, our Frame Annotation Module (FAM) ensures that we are able to generate high-quality and accurate captions for these videos. As shown in Fig. 10, without FAM, the generated caption is not able to describe the sudden appearance of the kangaroo’s right foot near its tail. Further, the caption generated without our FAM is also not able to describe the extra set of paws that appear suddenly from the legs of the cat. Thus, FAM plays an important role in curating high-quality and accurate video captions.
@@ Appendix D Implementation Details
We use the official code of each open-source Video-LMM for evaluation. Each of these codes are implemented in pytorch framework. We evaluate each one of them on a single NVIDIA A100 40GB GPU. For closed source Video-LMMs, we use their respective API for evaluation. We use GPT-4o OpenAI (2024) as our LMM to generate the captions and the final QA pairs in VANE-Bench. Next, we describe the prompts used in our Caption Generation Module (CGM), Question Answer Generation Module (QAGM), and in evaluating various Video-LMMs on VANE-Bench in the subsequent subsections.
@@ D.2 Question Answer Generation Module (QAGM)
Text Prompt: You are given a video input, which is generated by a state-of-the-art AI algorithm. Thus, these videos look very natural and almost realistic, but they are actually synthetic and generated by an AI algorithm. The videos may have some inconsistencies or anomalies present in them, which are generally localized to only a specific location in the video as identified by the green bounding boxes in the video. The rest of the video appears completely natural or realistic. This specific inconsistency may last for only a few frames of the video or may last for the entire video itself. The inconsistency or anomalies in the video are generally events and phenomena which is not observed in real-world and physical scenarios. You will also be given a caption as input that describes the video, along with the specific inconsistency present in the video. Based on the given video and caption input, your task is to formulate 3 diverse and misleading questions to test whether the multi-modal LLM model can correctly identify the options based on the inconsistencies present in the video or not. So, your generated questions should give the model few options to choose from to make its answer, and these options should be of high quality and also have misleading choices so that you can test deeper level of understanding of these multi-modal LLM models. Thus, the goal of these questions is to accurately assess the multi-modal LLM’s ability to accurately identify the inconsistencies present in the video. Generate questions that comprise both interrogative and declarative sentences, utilizing different language styles, and provide an explanation for each. Your response should be presented as a list of dictionary strings with keys ’Q’ for questions and ’A’ for the answer. Follow these rules while generating question and answers:
1. Do not provide answers in the question itself. For example, the ground-truth attribute or component that makes the video scene unusual should never be mentioned in the question itself.
@@
9. You must only follow the below output format and strictly must not output any other extra information or text. Your output format should be strictly as follows, without any additional information or text: